OpenStudy (abmon98):

the parametric equations of a curve are x=2theta-sin2theta y=2-cos2theta show that dy/dx=cot(theta) find the equation of the tangent to the curve at the point where theta=1/4 pi for the part of the curve where 0

OpenStudy (abmon98):

\[dy/dx=dy/d \theta*d \theta/dx\]

OpenStudy (abmon98):

\[2-\cos2\theta=2\sin2\theta \]

OpenStudy (anonymous):

\[\frac{dy}{d\theta}=2\sin2\theta\\ \frac{dx}{d\theta}=2-2\cos2\theta\] So \[\frac{dy}{dx}=\frac{2\sin2\theta}{2-2\cos2\theta}=\frac{\sin2\theta}{1-\cos2\theta}\] The tangent to the curve when \(\theta=\dfrac{\pi}{4}\) will pass through the point \(\left(x\left(\dfrac{\pi}{4}\right),y\left(\dfrac{\pi}{4}\right)\right)\) (the point on the curve at which \(\theta=\dfrac{\pi}{4}\)) and will have a slope of \(\dfrac{dy}{dx}\bigg|_{\theta=\pi/4}\) (the derivative evaluated at \(\theta=\dfrac{\pi}{4}\)). The tangent line will be parallel to the \(x\) axis (i.e. the line \(y=0\)) when \(\dfrac{dy}{dx}=0\), or simply when \(\dfrac{dy}{d\theta}=0\). Find the value of \(\theta\) between 0 and \(2\pi\) that makes this happen.

OpenStudy (abmon98):

Thank you so much :D